We installed the new UNe lamp into IGRINS and took some test images.
Ueejeong completed his software tests in the morning.
Greg and Hwi tested the IGRINS software and checked the systems for shipping.
In the afternoon the CTI 1050 Cold Head started to make noises and we found that the power cable had failed.
The cable was remade and power was resupplied to the CCR.
A list of software questions was sent to Ueejeong and he replied via email.
We started the day checking the software once more and then powered down IGRINS.
The Calibration Unit was loaded into its box and computers were packed.
We disconnected the electronics cabinet and helium lines and loaded IGIRNS onto the transportation cart.
Then the electronics cabinet and helium lines were reattached an IGRINS stayed cold overnight.
In the morning we completed packing the IGRINS accessories.
The compressor was disconnected from IGRINS and packed.
The glycol chiller was emptied of glycol and packed.
We started loading the truck at 3pm and were done around 4:15pm.
To avoid rain and traffic we left UT around 8pm and arrived at McDonald at 3:30am.
We started unloading the truck at 8:30am. Detailed procedures are on the Old Transportation Procedures page.
- TempTek: 1st floor, connected to glycol hoses
- Brooks He compressor: 3rd floor, connected to He hoses
- Electronic cabinet: dome, connected to read vacuum pressure, temperature of coldheads, grating etc
- ICC and DSC computers were installed in the control room, but not connected to the network.
Started vacuum pumping IGRINS at 9:30am.
- vacuum pump: started at 2.9x10^-2, temp ~200K
We let IGRINS passively warm while on the pump.
Started cooling at 8:30am.
- vacuum at 3.2x10^-4, temperature 250K.
Closed the valve and turned off the pump at 10:30am.
- vacuum at 2.5x10^-7, temperature 240K.
Connected the computers to the network at 9:30am.
- Setup ICC & DSC
- Network is setup with a McD DNS server:
- NOTE: DSC (192.168.2.1) is a router, ICC (192.168.2.2) is connected to ethernet through DSC.
Connected the detectors and fiber optic cable to IGRINS and initialized the detectors.
Ueejeong started software testing.
IGRINS Installed onto telescope.
First night of observing.
SVC socket issue. I'll attach Ueejeong's email later.
The vacuum pressure was slightly higher (10^-6) than yesterday (10^-8).
The He compressor was not getting any chilled water supply so I checked the Temptek chiller and the glycol level was low.
Contacted Coyne and found that the main reservoir's glycol lost ~20 gallon and temptek lost ~4 gallon.
Filled with tap water and ice to cool it down quickly and the temperature of the chiller went down to 53.
The supply/return lines were not as cool as I remembered...